Joellen Company’s revenues are $300,000 and invested capital is $240,000. Expenses are currently 60% of sales. Joellen Company’s current return on investment is

Group of answer choices

100 %

80 %

50 %

None of these answers is correct
Answer #1

Answer is option C

c. 50%

Return on investment = net income / invested capital

Net income = revenues – expenses = 300000-(300000*60%) = $120000

Return on investment = 120000/240000 = 50%